OpenStudy (anonymous):

will fan and medal Use the correct order of operations to solve the problem below. 3 × 9 + 6 ÷ 2 - (3 + 2) 29 25 35 31

OpenStudy (texaschic101):

3 * 9 + 6/2 - (3 + 2) parenthesis first.. 3 * 9 + 6/2 - 5 now do multiplication/division from left to right 27 + 3 - 5 now do addition/subtraction from left to right 30 - 5 answer is : 25

OpenStudy (anonymous):

thanks can you help with more

OpenStudy (texaschic101):

sure :)

OpenStudy (anonymous):

Use the correct order of operations to solve the problem below. 6 + 18 ÷ 3 8 11 12 27

OpenStudy (texaschic101):

6 + 18/3 --- do division first 6 + 6 -- now add 12

OpenStudy (texaschic101):

order of operations : PEMDAS -- in that order P == parenthesis (first) E == exponents (if there is any) M/D == multiplication/division from left to right A/S == addition/subtraction from left to right

OpenStudy (anonymous):

Evaluate the expression below at x = 4. 36 12 48 1,296

OpenStudy (texaschic101):

can't see the expression

OpenStudy (texaschic101):

(3x - 6) ^ (x-2)...when x = 4 (3(4) - 6) ^ (4-2) = (12 - 6) ^ 2 = 6^2 = 36

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

Evaluate the expression below at x = 9. 7 54 774 14